2x-13/5-x-3/11=x-9/5+1




Answer by stanbon(75887)

(2x-13)/5 - (x-3)/11 = (x-9)/5 + 1
---------------------
Multiply thru by 55 to get:
11(2x-13) - 5(x-3) = 11(x-9) + 11
------
22x - 143 - 5x + 15 = 11x - 99 + 11
-----------------------------------------
17x - 128 = 11x - 88
------
6x = 40
----
x = 6 2/3
===================
Cheers,
Stan H.